OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 3x2 − 3x − 60.

OpenStudy (anonymous):

so how do i factor it or wutever?

OpenStudy (anonymous):

3x2 − 3x − 60=3(x^2 -x -20) =3(x^2 -5x+4x -20) =3(x(x-5)+4(x-5))=3(x-5)(x+4)